# Community-Centered Spatial Intelligence for Climate Adaptation at Nova Scotia’s Eastern Shore

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: 1st Workshop on Spatial Intelligence for Smart and Connected Communities; Nov 03–06, 2025; Minneapolis, MN, USAisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Spatial-temporal systemsccs: Applied computing Law, social and behavioral sciences
@@
Beach Profiles Monthly surveys using a Single-User Beach Profiler (SUBP) (1). To quantitatively track coastal erosion versus climatic seasonal changes. This provides communities with hard evidence to support planning and adaptation decisions.
@@ 1. Introduction
Long before it was included in official maps, the Eastern Shore of Nova Scotia was a living territory, the ancestral and unceded home of the Mi’kmaq people, known as Sipekne’katik and Eskikewa’kik (2; 3). For millennia, they lived in harmony with the tides, the seasons, and the abundant life of the coast (4). European settlement brought changes, with Canso emerging as a fishing station in the early seventeenth century (5). Over the next three hundred years, a unique culture took root. Small, hardy villages sprang up in sheltered coves and harbors, their existence dictated by the ocean (6).
Life on the Eastern Shore was defined by a culture of profound self-reliance (7; 8). Communities were built around inshore fishing, boat building, and logging (9). The communal fish shed and the lobster cannery were the economic and social hearts of these villages (10). Due to the surrounding environment, people of the shore developed an intimate, almost hereditary knowledge of their environment (11). They could read the weather in the clouds, understand the temperament of the tides, and navigate the treacherous and foggy archipelago of over a thousand wild islands that guard the coast (12). Their history is deeply ingrained in the landscape and preserved through folklore, sea shanties, and tales of shipwrecks.
Today, Nova Scotia’s Eastern Shore faces an economic shift that has led to the decline of traditional industries, prompting many younger people to seek opportunities elsewhere, leaving behind an aging population (13). At the same time, climate change presents a more insidious peril (14). Rising sea levels, intensifying storms, and coastal erosion are already a reality that threaten homes, wharves, and the very land (15). The project described in this short paper was created from this context, not as an external intervention, but as a collaborative effort to fortify this historic culture against the coming changes. We aim to bridge the profound, intuitive wisdom of the community with the analytical power of modern science, ensuring the future of the Eastern Shore remains rooted in its past.
@@ 2. Collaborative Coastal Monitoring
At the center of this initiative is the Eastern Shore Citizen Science Coastal Monitoring Network (ESCOM). This community-based program emerged not from academic laboratories but from the coastal landscapes it now monitors (16). Originating through a 2024 pilot led by The Rev. Marian Lucas-Jeferies and Dr. Camilo M. Botero, with the underlying intent of connecting science and community (17), the network has since matured both physically and digitally (18) into a structured and locally embedded effort dedicated to advancing coastal and climate literacy and resilience.
The network’s approach is grounded in a philosophy of direct community empowerment (19). Rather than positioning residents as passive participants, the project invites them to become active collaborators in the scientific process. Participants are encouraged to define their research questions, apply systematic methods to collect data, and contribute substantively to shared understandings of climate risks. In doing so, the network not only generates valuable observational datasets but also reinforces the principle that local knowledge holders are essential players in climate governance.
@@
Local institutions are crucial in this ecosystem (16; 20). Nonprofit organizations such as The Deanery Project and the Port Bickerton and Area Planning Association, alongside the Anglican Parish of Port Dufferin and the five schools in the area, provide spaces for recruitment, training, and sustained dialogue. These community anchors enable a model of knowledge co-production that deliberately integrates historical insight with contemporary technical skills. Elders offer decades of observational experience and cultural memory, while younger participants contribute digital expertise and novel analytical perspectives. Through this intergenerational collaboration, the network cultivates a living record of coastal change, one that is deeply informed by place, community, and shared responsibility.
@@ 3.1. Beach Profile Surveys
Every month, volunteer teams conduct beach profile surveys. They use a Single-User Beach Profiler (SUBP) (1), constructed from everyday materials such as curtain rods, wing nuts, and a pocket level, to measure the cross-sectional shape of their local beach. This low-tech application of the SUBP method is particularly precise and can detect subtle changes in sand elevation (21). This ability enables the community to distinguish between normal seasonal fluctuations and long-term erosion along the coastline. The collected data provides a solid, empirical baseline for assessing the impact of storms or human interventions on nearby coastal communities.

### v2

@@ Abstract.
journalyear: 2025copyright: ccconference: The 1st ACM SIGSPATIAL International Workshop on Spatial Intelligence for Smart and Connected Communities; November 3–6, 2025; Minneapolis, MN, USAbooktitle: The 1st ACM SIGSPATIAL International Workshop on Spatial Intelligence for Smart and Connected Communities (SpatialConnect ’25), November 3–6, 2025, Minneapolis, MN, USAdoi: 10.1145/3764924.3770892isbn: 979-8-4007-2187-8/2025/11ccs: Information systems Spatial-temporal systemsccs: Applied computing Law, social and behavioral sciences
@@
Beach Profiles Monthly surveys using a Single-User Beach Profiler (SUBP) (Puleo* et al., 2008). To quantitatively track coastal erosion versus climatic seasonal changes. This provides communities with hard evidence to support planning and adaptation decisions.
@@ 1. Introduction
Long before it was included in official maps, the Eastern Shore of Nova Scotia was a living territory, the ancestral and unceded home of the Mi’kmaq people, known as Sipekne’katik and Eskikewa’kik (Sipekne’katik First Nation, ; Lewis and Sable, 2025). For millennia, they lived in harmony with the tides, the seasons, and the abundant life of the coast (Pastore, 1998). European settlement brought changes, with Canso emerging as a fishing station in the early seventeenth century (Parks Canada, ). Over the next three hundred years, a unique culture took root. Small, hardy villages sprang up in sheltered coves and harbors, their existence dictated by the ocean (Fisherman’s Life Museum, 2018).
Life on the Eastern Shore was defined by a culture of profound self-reliance (Munroe, 1996; Society, 2001). Communities were built around inshore fishing, boat building, and logging (Project, 2018). The communal fish shed and the lobster cannery were the economic and social hearts of these villages (Society, 2019). Due to the surrounding environment, people of the shore developed an intimate, almost hereditary knowledge of their environment (Marit Berger et al., 2024). They could read the weather in the clouds, understand the temperament of the tides, and navigate the treacherous and foggy archipelago of over a thousand wild islands that guard the coast (Trust, 2025). Their history is deeply ingrained in the landscape and preserved through folklore, sea shanties, and tales of shipwrecks.
Today, Nova Scotia’s Eastern Shore faces an economic shift that has led to the decline of traditional industries, prompting many younger people to seek opportunities elsewhere, leaving behind an aging population (Bradley, 2018). At the same time, climate change presents a more insidious peril (of Natural Resources and Renewables, 2021). Rising sea levels, intensifying storms, and coastal erosion are already a reality that threaten homes, wharves, and the very land (College, 2024). The project described in this short paper was created from this context, not as an external intervention, but as a collaborative effort to fortify this historic culture against the coming changes. We aim to bridge the profound, intuitive wisdom of the community with the analytical power of modern science, ensuring the future of the Eastern Shore remains rooted in its past.
@@ 2. Collaborative Coastal Monitoring
At the center of this initiative is the Eastern Shore Citizen Science Coastal Monitoring Network (ESCOM). In this paper, citizen scientists are non-professional volunteers (primarily seniors) who collaborate with researchers to help define questions, collect and analyze data, and interpret results, thereby generating new scientific knowledge (Law, ). The program emerged not from academic laboratories but from the coastal landscapes it now monitors (Attree, 2025). Originating in a 2024 pilot led by The Rev. Marian Lucas-Jeferies and Dr. Camilo M. Botero with the intent of connecting science and community (Eberhardt et al., 2022), the network has since matured both physically and digitally (Lucrezi, 2021) into a structured and locally embedded effort dedicated to advancing coastal and climate literacy and resilience.
The network’s approach is grounded in a philosophy of direct community empowerment (Bell, 2025). Rather than positioning residents as passive participants, the project invites them to become active collaborators in the scientific process. Participants are encouraged to define their research questions, apply systematic methods to collect data, and contribute substantively to shared understandings of climate risks. In doing so, the network not only generates valuable observational datasets but also reinforces the principle that local knowledge holders are essential players in climate governance.
@@
Local institutions are crucial in this ecosystem (Attree, 2025; Jefferies, 2024). Nonprofit organizations such as The Deanery Project and the Port Bickerton and Area Planning Association, alongside the Anglican Parish of Port Dufferin and the five schools in the area, provide spaces for recruitment, training, and sustained dialogue. These community anchors enable a model of knowledge co-production that deliberately integrates historical insight with contemporary technical skills. Elders offer decades of observational experience and cultural memory, while younger participants contribute digital expertise and novel analytical perspectives. Through this intergenerational collaboration, the network cultivates a living record of coastal change, one that is deeply informed by place, community, and shared responsibility.
@@ 3.1. Beach Profile Surveys
Every month, volunteer teams conduct beach profile surveys. They use a Single-User Beach Profiler (SUBP) (Puleo* et al., 2008), constructed from everyday materials such as curtain rods, wing nuts, and a pocket level, to measure the cross-sectional shape of their local beach. This low-tech application of the SUBP method is particularly precise and can detect subtle changes in sand elevation (Torres et al., 2023). This ability enables the community to distinguish between normal seasonal fluctuations and long-term erosion along the coastline. The collected data provides a solid, empirical baseline for assessing the impact of storms or human interventions on nearby coastal communities.